# CLASSP: a Biologically-Inspired Approach to Continual Learning through Adjustment Suppression and Sparsity Promotion

### v1

@@ Abstract
This paper introduces a new biologically-inspired training method named Continual Learning through Adjustment Suppression and Sparsity Promotion (CLASSP). CLASSP is based on two main principles observed in neuroscience, particularly in the context of synaptic transmission and Long-Term Potentiation (LTP). The first principle is a decay rate over the weight adjustment, which is implemented as a generalization of the AdaGrad optimization algorithm. This means that weights that have received many updates should have lower learning rates as they likely encode important information about previously seen data. However, this principle results in a diffuse distribution of updates throughout the model, as it promotes updates for weights that haven’t been previously updated, while a sparse update distribution is preferred to leave weights unassigned for future tasks. Therefore, the second principle introduces a threshold on the loss gradient. This promotes sparse learning by updating a weight only if the loss gradient with respect to that weight is above a certain threshold, i.e. only updating weights with a significant impact on the current loss. Both principles reflect phenomena observed in LTP, where a threshold effect and a gradual saturation of potentiation have been observed. CLASSP is implemented in a Python/PyTorch class, making it applicable to any model. When compared with Elastic Weight Consolidation (EWC) using Computer Vision datasets, CLASSP demonstrates superior performance in terms of accuracy and memory footprint.
@@ 1 Introduction
The paper is organized as follows. Section 2 briefly reports the state-of-the-art in Continual Learning and Catastrophic Forgetting, Section 3 justifies and explains the proposed training algorithm. Section 4 shows the effectiveness of CLASSP, through an ablation study and its comparison with AdaGrad and Elastic Weight Consolidation (EWC) methods using Computer Vision datasets, and conclusions are presented in Section 5.
@@ 3 CLASSP
0:$params:$ learning rate $\alpha$, thresholdthreshold, power $p$, apply_decayapply\_decay and $\epsilon$
0:lossloss
1: Initialize CLASSP with $\alpha$, thresholdthreshold, power $p$, apply_decayapply\_decay and $\epsilon$
@@
3:  Calculate lossloss with autograd
@@
8:     Initialize grad_sumgrad\_sum for $w$ if not already done
@@
12:      if apply_decayapply\_decay is True then
@@ 4 Experiments
This section presents the results of our experiments to evaluate the performance of CLASSP in the continual learning setting using two popular computer vision datasets, MNIST and Fashion MNIST (FMNIST). For all experiments, training is first performed for 4 epochs on the MNIST dataset and then resumes for one epoch on the FMNIST dataset, allowing us to assess the degree of forgetting on the first dataset. Figure 1 shows the difference between the two datasets.
@@
The model architecture leverages a Convolutional Neural Network (CNN) implemented in PyTorch. The CNN comprises two convolutional layers, each with 3x3 filters, interspersed with dropout layers for regularization, and two fully-connected layers for final classification. The first convolutional layer applies 32 filters, followed by a ReLU activation function. The second convolutional layer utilizes 64 filters and is followed by a ReLU activation and a max pooling operation. The loss function is cross-entropy for all experiments except for EWC.
The results summarized in Table 1 provide a comparative analysis of various training algorithms on the sequence of MNIST and FMNIST datasets. A series of 10 experiments per algorithm were conducted, from which the mean accuracy and standard deviation were derived. To isolate the issue of forgetting from the model’s capacity to generalize, the same datasets were used for both training and evaluation.
The best configuration for CLASSP was established with learning rate of 0.2, power $p=1$ and thresholdthreshold of 0.5, which was applied only to the first dataset, to leave unassigned weights for the second tasks. The decay rate was applied exclusively to the second dataset, i.e. the optimizer computes the sum of the gradients while processing the first dataset, but applies the decay rate only to the second dataset. This process is detailed in Lines 10-15 of Algorithm 1.
@@
These experimental results demonstrate the effectiveness of CLASSP in mitigating catastrophic forgetting during continual learning. The best configuration for CLASSP achieved a mean accuracy of $91.05\%$ on MNIST and $85.53\%$ on FMNIST, outperforming the baseline algorithms. Notably, the introduction of a threshold in CLASSP significantly improved performance on MNIST compared to the version without a threshold, indicating the importance of this parameter in the learning process.
@@ 5 Conclusions
Compared to existing methods like EWC, CLASSP offers advantages in terms of memory footprint and performance. It only requires storing weight-specific scaling factors, as opposed to EWC’s need to buffer all past weights and elements of the Fisher Information Matrix. Experiments on MNIST and Fashion-MNIST datasets demonstrate CLASSP’s effectiveness, achieving higher average accuracy compared to AdaGrad, SGD and EWC, validating CLASSP as a promising method for continual learning, with potential applications in various domains where learning from sequential or streaming data is crucial.
As future work, I see promise in applying CLASSP to more complex datasets and tasks beyond computer vision. Investigating the impact of scale factor quantization for further memory efficiency is also of interest. Finally, it would be valuable to delve deeper into the theoretical aspects of CLASSP’s convergence properties. This could, for example, start by examining Theorem 6 of [9], taking into account the added complexity introduced by the thresholding mechanism, to gain a better understanding of the interplay between the adaptive learning rate and the thresholding mechanism.

### v2

@@ Abstract
This paper introduces a new biologically-inspired training method named Continual Learning through Adjustment Suppression and Sparsity Promotion (CLASSP). CLASSP is based on two main principles observed in neuroscience, particularly in the context of synaptic transmission and Long-Term Potentiation (LTP). The first principle is a decay rate over the weight adjustment, which is implemented as a generalization of the AdaGrad optimization algorithm. This means that weights that have received many updates should have lower learning rates as they likely encode important information about previously seen data. However, this principle results in a diffuse distribution of updates throughout the model, as it promotes updates for weights that haven’t been previously updated, while a sparse update distribution is preferred to leave weights unassigned for future tasks. Therefore, the second principle introduces a threshold on the loss gradient. This promotes sparse learning by updating a weight only if the loss gradient with respect to that weight is above a certain threshold, i.e. only updating weights with a significant impact on the current loss. Both principles reflect phenomena observed in LTP, where a threshold effect and a gradual saturation of potentiation have been observed. CLASSP is implemented in a Python/PyTorch class, making it applicable to any model. When compared with Elastic Weight Consolidation (EWC) using computer vision and sentiment analysis datasets, CLASSP demonstrates superior performance in terms of accuracy and memory footprint.
@@ 1 Introduction
The paper is organized as follows. Section 2 briefly reports the state-of-the-art in Continual Learning and Catastrophic Forgetting, Section 3 justifies and explains the proposed training algorithm. Section 4 shows the effectiveness of CLASSP through an ablation study and its comparison with AdaGrad, Adam, SGD and Elastic Weight Consolidation (EWC) methods using computer vision and sentiment analysis datasets. Conclusions are presented in Section 5.
@@ 3 CLASSP
0:  $params:$ learning rate $\alpha$, $threshold$, power $p$, $apply\_decay$ and $\epsilon$
0:  $loss$
1:  Initialize CLASSP with $\alpha$, $threshold$, power $p$, $apply\_decay$ and $\epsilon$
@@
3:     Calculate $loss$ with autograd
@@
8:              Initialize $grad\_sum$ for $w$ if not already done
@@
12:                 if $apply\_decay$ is True then
@@ 4 Experiments
This section presents the results of our experiments to evaluate the performance of CLASSP in the continual learning setting. The first set of experiments uses two popular computer vision datasets: MNIST [9] and Fashion MNIST (FMNIST) [10]. The second set of experiments uses two sentiment analysis datasets: Financial Phrasebank (FPB) [11] and IMDB [12]. To isolate the issue of forgetting from the model’s capacity to generalize, the same datasets were used for both training and evaluation.
4.1 Computer Vision
For all computer vision experiments, the model is trained for 4 epochs on MNIST, followed by 1 epoch on FMNIST. This setup allows assessment of how well the model retains knowledge of MNIST after learning FMNIST. Figure 1 shows the difference between the two datasets.
@@ 4.1 Computer Vision
The model architecture leverages a Convolutional Neural Network (CNN) implemented in PyTorch [^2] See the script for the experiment with CLASSP in computer vision at https://github.com/oswaldoludwig/CLASSP/blob/main/experiment_CV.py. The CNN comprises two convolutional layers, each with 3x3 filters, interspersed with dropout layers for regularization, and two fully-connected layers for final classification. The first convolutional layer applies 32 filters, followed by a ReLU activation function. The second convolutional layer utilizes 64 filters and is followed by a ReLU activation and a max pooling operation. The loss function is cross-entropy for all experiments except for EWC [^3] See the EWC script in computer vision at https://github.com/oswaldoludwig/CLASSP/blob/main/experiment_EWC.py.
The results summarized in Table 1 provide a comparative analysis of various training algorithms on the sequence of MNIST and FMNIST datasets. A series of 10 experiments per algorithm were conducted, from which the mean accuracy and standard deviation were derived.
The best configuration for CLASSP was established with learning rate of 0.2, power $p=1$ and $threshold$ of 0.5, which was applied only to the first dataset, to leave unassigned weights for the second tasks. The decay rate was applied exclusively to the second dataset, i.e. the optimizer computes the sum of the gradients while processing the first dataset, but applies the decay rate only to the second dataset. This process is detailed in Lines 10-15 of Algorithm 1.
@@
4.2 Text Classification
In all sentiment analysis experiments, the FPB dataset is used for initial training, followed by further training on the IMDB dataset [^4] See the script for the experiment with CLASSP in sentiment analysis at https://github.com/oswaldoludwig/CLASSP/blob/main/experiment_SA.py. Both datasets involve sentiment analysis, but differ in context: FPB focuses on financial news articles, aiming to predict the impact on stock prices, while IMDB consists of movie reviews, providing sentiment analysis for the entertainment domain.
The experiments leverage a Transformer Encoder for text classification. The model first embeds the input text using a 600-dimensional embedding layer. Dropout with a rate of 0.1 and layer normalization are then applied. The processed signal is then fed into a single-layer Transformer encoder with 30 attention heads. Finally, the output is normalized and fed into a linear layer with three units, corresponding to the three sentiment classes (positive, neutral, negative) based on the FPB annotation scheme.
To prevent overfitting on the IMDB dataset, which could worsen forgetting of the FPB dataset, we employ a threshold on the loss function during IMDB training. This threshold, which is a function of the optimizer, maintains IMDB accuracy at approximately $85\%$ ensuring a fair comparison. While all optimizers achieved over $99\%$ accuracy on FPB, CLASSP exhibited superior performance in mitigating forgetting. The forgetting rate observed with CLASSP ($33.34\%$) was significantly lower compared to Adam ($53.09\%$) or even EWC [^5] See the EWC script in sentiment analysis at https://github.com/oswaldoludwig/CLASSP/blob/main/experiment_SA_EWC.py ($47.15\%$).
To understand the contribution of different components in CLASSP, Table 2 presents the results of an ablation study, alongside baseline performance of SGD, Adam, and EWC.
Table 2: Comparative analysis of different training algorithms on the dataset sequence FPB + IMDB: mean accuracy and standard deviation calculated over 10 experiments.
training algorithm dataset mean std
CLASSP FPB 66.04 3.85
$\alpha=0.025$, threshold=0.99, p=1 IMDB 85.74 1.30
CLASSP without threshold FPB 64.81 6.63
$\alpha=0.025$, threshold=0.0, p=1 IMDB 85.48 1.38
CLASSP (AdaGrad instance) FPB 52.95 1.97
$\alpha=0.025$, threshold=0.0, p=2 IMDB 85.01 2.53
Vanilla SGD FPB 47.24 5.57
IMDB 85.67 1.72
Adam FPB 46.63 3.72
IMDB 85.66 1.99
EWC FPB 52.46 1.49
IMDB 85.67 1.15
Once again $p=1$ is the optimal power. The power $p$ in both Adagrad and CLASSP algorithms determines how the past gradients are accumulated and used to adjust the learning rate. In Adagrad, $p=2$ means that larger gradients have a stronger influence in decreasing the learning rate, as they are squared before being accumulated. This can lead to a rapid decrease in the learning rate, especially if large gradients are encountered early in training, as usual. On the other hand, CLASSP uses the sum of absolute values of past gradients, i.e. $p=1$. This means that all gradients, regardless of their magnitude, contribute linearly to the accumulated sum, leading to a more balanced and slower decrease of the learning rate, as it is less sensitive to large gradients.
In summary, these experiments encompassing both computer vision and sentiment analysis tasks demonstrate the effectiveness of CLASSP in mitigating catastrophic forgetting during continual learning.
@@ 5 Conclusions
Compared to existing methods like EWC, CLASSP offers advantages in terms of memory footprint and performance. It only requires storing weight-specific scaling factors, as opposed to EWC’s need to buffer all past weights and elements of the Fisher Information Matrix. Experiments on MNIST, Fashion-MNIST, IMDB and Financial Phrasebank datasets demonstrate CLASSP’s effectiveness, achieving higher average accuracy compared to AdaGrad, SGD, Adam and EWC, validating CLASSP as a promising method for continual learning, with potential applications in various domains where learning from sequential or streaming data is crucial.
As future work, I see promise in applying CLASSP to more complex datasets and tasks beyond computer vision and text classification. Investigating the impact of scale factor quantization for further memory efficiency is also of interest. Finally, it would be valuable to delve deeper into the theoretical aspects of CLASSP’s convergence properties. This could, for example, start by examining Theorem 6 of [13], taking into account the added complexity introduced by the thresholding mechanism, to gain a better understanding of the interplay between the adaptive learning rate and the thresholding mechanism.